How many different anagrams (including nonsensical words) can be made from the letters in the word STATISTICS, using all the letters?
Answer #1

Solution:

STATISTICS contains 10 letters, where we have 3S, 3Ts, 2Is, and 1 letter of the other letters

N= 10

S = 3

I = 2

T = 3

A = 1

C = 1

Number of anargrams = (10!)/((3! *2! * 3!* 1! *1!) = 3628800/(6*2*6*1*1) = 50400

So there is 50400 different anargrams.